# Time-To-Inconsistency: A Survival Analysis of Large Language Model Robustness to Adversarial Attacks

### v2

@@ 1 Introduction
Large Language Models (LLMs) have demonstrated remarkable capabilities across diverse tasks brown2020language; chowdhery2022palm; touvron2023llama, yet their deployment in high-stakes applications necessitates rigorous evaluation of their consistency under adversarial conditions hendrycks2021measuring; lin2022truthfulqa. While existing evaluation frameworks primarily assess single-turn performance liang2022holistic; eval-harness, real-world interactions involve sustained multi-turn conversations where models must maintain consistency despite evolving contexts and adversarial pressure shuster2022blenderbot; bai2022training.
Current evaluation paradigms exhibit fundamental limitations in capturing the temporal dynamics of conversational robustness kiela2021dynabench; ribeiro2020beyond. Standard benchmarks measure performance in isolated turns, and even multi-turn protocols are usually summarized by static aggregate scores. These views obscure how errors emerge and propagate over time: they cannot distinguish between a model that fails immediately under mild adversarial pressure and one that remains stable for many turns before eventually degrading. Phenomena such as sycophancy—where models readily abandon correct responses under minimal user challenges sharma2023towards; turpin2023language—illustrate that the trajectory of a conversation matters just as much as its final outcome.
Consider a medical assistant that initially provides accurate information but gradually shifts recommendations under persistent questioning singhal2022large; nori2023capabilities, or a system that maintains precision for straightforward queries yet fails catastrophically only after a specific pattern of semantic drift and adversarial prompts zou2023universal; wei2024jailbroken. In both cases, what makes these failures concerning is not just that they occur, but when they occur and how they are precipitated by the dialogue history. From a safety and reliability perspective, we need tools that can answer questions such as: How quickly do errors emerge under different adversarial strategies? Which kinds of semantic shifts most sharply increase the risk of failure? And can we identify conversations that are on a “high-risk” trajectory before an error actually occurs?
@@
We address these questions by framing multi-turn robustness as a time-to-event problem and analyzing the time-to-inconsistency of an LLM within an adversarial conversation. The event is the first incorrect answer under a strict consistency criterion; time is measured in discrete turns; and conversations that remain correct within an 8-turn horizon are treated as right-censored. Survival analysis cox1972regression; kalbfleisch2011statistical naturally fits this setting: it separates whether a conversation fails from when it fails, handles censored dialogues without ad hoc labels, and provides turn-wise hazard functions that track how failure risk evolves over the dialogue. Because survival models support time-varying covariates, they also let us link evolving conversational signals directly to changes in risk.
We instantiate this framework on the MT-Consistency benchmark li-etal-2025-firm, analyzing 36,951 turns from 9 state-of-the-art LLMs using Cox proportional hazards, Accelerated Failure Time (AFT), and Random Survival Forest (RSF) models to understand which assumptions best capture multi-turn failure dynamics. This work makes three main contributions:
@@ 2.1 Multi-Turn Degradation and Evaluation in LLMs
Recent research consistently demonstrates that large language models (LLMs) exhibit significant performance degradation during multi-turn interactions compared to single-turn tasks laban2025llms; li2025beyond. This degradation manifests primarily as increased inconsistency and variance across conversational turns, arising from premature conclusions and overly confident reliance on incorrect intermediate responses laban2025llms. To systematically measure such inconsistencies, several specialized benchmarks have been developed. Early frameworks such as MT-Bench zheng2023judging primarily evaluated two-turn interactions, while subsequent efforts like MT-Bench-101 bai2024mt extended these evaluations to more extensive dialogue scenarios, highlighting uneven multi-turn performance even in advanced chat-tuned models. Complementarily, MT-Eval kwan2024mt introduced controlled experiments to explicitly contrast single-turn and multi-turn performance, identifying error propagation and distant contextual dependencies as critical contributors to performance decline. Additionally, benchmarks like MultiChallenge sirdeshmukh2025multichallenge emphasize realistic conversational complexities, exposing significant limitations in current models’ ability to manage ambiguous instructions and context shifts across turns.
@@ 2.2 Consistency and Sycophantic Behavior
Focused examinations into specific multi-turn failure modes have uncovered critical phenomena such as ”sycophantic drift,” where models alter correct answers in response to user pushback or misleading follow-ups. The FlipFlop Experiment by laban2023you empirically demonstrated this vulnerability, observing frequent reversals from correct to incorrect answers under trivial user challenges. To quantify and mitigate this issue, li-etal-2025-firm introduced the Position-Weighted Consistency (PWC) metric, penalizing early-stage inconsistencies due to their detrimental impact on user trust. Their Confidence-Aware Response Generation (CARG) method notably improved multi-turn consistency by leveraging the model’s internal confidence signals. Our hazard-modeling approach complements these findings by statistically characterizing the increasing risk of response inconsistency over dialogue turns.
@@ 2.3 Survival Analysis and Sequential Modeling
Survival analysis techniques, traditionally employed to model time-to-event data cox1972regression; kalbfleisch2011statistical, has recently been applied to conversational settings, e.g., to predict dialogue termination or disruptions de-kock-vlachos-2021-survival; maystre2022temporally. These works, however, focus on user- or session-level outcomes and do not address the internal consistency of LLM responses under adversarial pressure. In contrast, we model time-to-inconsistency: the first incorrect answer in a multi-turn adversarial dialogue. We combine Cox proportional hazards, Accelerated Failure Time, and Random Survival Forest models and link their behavior to semantic drift covariates, enabling a nuanced statistical characterization of error accumulation and offering novel insights into dialogue reliability dynamics previously observed only empirically.
@@ 3.2 Predictive Feature Engineering
We encode each user prompt with a sentence-transformer model reimers2019sentencebert:
@@ 3.3 Survival Modeling Framework
As a flexible non-parametric baseline, we employ Random Survival Forests (RSF) ishwaran2008random, which fit an ensemble of survival trees on bootstrap samples. At each split, candidate covariates are sampled at random and chosen to maximize a survival impurity reduction (log-rank statistic). Each terminal node yields a Nelson–Aalen estimate of the cumulative hazard; the forest prediction for conversation $i$ is obtained by averaging cumulative hazards across trees and converting to survival probabilities. RSF can capture nonlinearities and high-order interactions between drift features and model identity without explicit parametric assumptions.
@@ 4.1 Data
We conduct our study on the MT-Consistency li-etal-2025-firm, which systematically probes LLM consistency under adversarial multi-turn interactions. Each conversation is built from a base question followed by up to 8 adversarial follow-ups; we adopt this 8-turn horizon in all experiments.
@@ B.1 Complete Subject-to-Cluster Mappings
computer science, computer_security,
@@
General Knowledge (2 subjects) truthful_qa, common sense

### v3

@@ 1 Introduction
Large Language Models (LLMs) have demonstrated remarkable capabilities across diverse tasks Brown et al. (2020); Chowdhery et al. (2023); Touvron et al. (2023), yet their deployment in high-stakes applications necessitates rigorous evaluation of their consistency under adversarial conditions Hendrycks et al. (2020); Lin et al. (2022). While existing evaluation frameworks primarily assess single-turn performance Liang et al. (2022); Gao et al. (2024), real-world interactions involve sustained multi-turn conversations where models must maintain consistency despite evolving contexts and adversarial pressure Shuster et al. (2022); Bai et al. (2022).
Current evaluation paradigms exhibit fundamental limitations in capturing the temporal dynamics of conversational robustness Kiela et al. (2021); Ribeiro et al. (2020). Standard benchmarks measure performance in isolated turns, and even multi-turn protocols are usually summarized by static aggregate scores. These views obscure how errors emerge and propagate over time: they cannot distinguish between a model that fails immediately under mild adversarial pressure and one that remains stable for many turns before eventually degrading. Phenomena such as sycophancy—where models readily abandon correct responses under minimal user challenges Sharma et al. (2023); Turpin et al. (2023)—illustrate that the trajectory of a conversation matters just as much as its final outcome.
Consider a medical assistant that initially provides accurate information but gradually shifts recommendations under persistent questioning Singhal et al. (2023); Nori et al. (2023), or a system that maintains precision for straightforward queries yet fails catastrophically only after a specific pattern of semantic drift and adversarial prompts Zou et al. (2023); Wei et al. (2023). In both cases, what makes these failures concerning is not just that they occur, but when they occur and how they are precipitated by the dialogue history. From a safety and reliability perspective, we need tools that can answer questions such as: How quickly do errors emerge under different adversarial strategies? Which kinds of semantic shifts most sharply increase the risk of failure? And can we identify conversations that are on a “high-risk” trajectory before an error actually occurs?
@@
We address these questions by framing multi-turn robustness as a time-to-event problem and analyzing the time-to-inconsistency of an LLM within an adversarial conversation. The event is the first incorrect answer under a strict consistency criterion; time is measured in discrete turns; and conversations that remain correct within an 8-turn horizon are treated as right-censored. Survival analysis Cox (1972); Kalbfleisch and Prentice (2002) naturally fits this setting: it separates whether a conversation fails from when it fails, handles censored dialogues without ad hoc labels, and provides turn-wise hazard functions that track how failure risk evolves over the dialogue. Because survival models support time-varying covariates, they also let us link evolving conversational signals directly to changes in risk.
We instantiate this framework on the MT-Consistency benchmark Li et al. (2025a), analyzing 36,951 turns from 9 state-of-the-art LLMs using Cox proportional hazards, Accelerated Failure Time (AFT), and Random Survival Forest (RSF) models to understand which assumptions best capture multi-turn failure dynamics. This work makes three main contributions:
@@ 2.1 Multi-Turn Degradation and Evaluation in LLMs
Recent research consistently demonstrates that large language models (LLMs) exhibit significant performance degradation during multi-turn interactions compared to single-turn tasks Laban et al. (2025); Li et al. (2025b). This degradation manifests primarily as increased inconsistency and variance across conversational turns, arising from premature conclusions and overly confident reliance on incorrect intermediate responses Laban et al. (2025). To systematically measure such inconsistencies, several specialized benchmarks have been developed. Early frameworks such as MT-Bench Zheng et al. (2023) primarily evaluated two-turn interactions, while subsequent efforts like MT-Bench-101 Bai et al. (2024) extended these evaluations to more extensive dialogue scenarios, highlighting uneven multi-turn performance even in advanced chat-tuned models. Complementarily, MT-Eval Kwan et al. (2024) introduced controlled experiments to explicitly contrast single-turn and multi-turn performance, identifying error propagation and distant contextual dependencies as critical contributors to performance decline. Additionally, benchmarks like MultiChallenge Deshpande et al. (2025) emphasize realistic conversational complexities, exposing significant limitations in current models’ ability to manage ambiguous instructions and context shifts across turns.
@@ 2.2 Consistency and Sycophantic Behavior
Focused examinations into specific multi-turn failure modes have uncovered critical phenomena such as ”sycophantic drift,” where models alter correct answers in response to user pushback or misleading follow-ups. The FlipFlop Experiment by Laban et al. (2023) empirically demonstrated this vulnerability, observing frequent reversals from correct to incorrect answers under trivial user challenges. To quantify and mitigate this issue, Li et al. (2025a) introduced the Position-Weighted Consistency (PWC) metric, penalizing early-stage inconsistencies due to their detrimental impact on user trust. Their Confidence-Aware Response Generation (CARG) method notably improved multi-turn consistency by leveraging the model’s internal confidence signals. Our hazard-modeling approach complements these findings by statistically characterizing the increasing risk of response inconsistency over dialogue turns.
@@ 2.3 Survival Analysis and Sequential Modeling
Survival analysis techniques, traditionally employed to model time-to-event data Cox (1972); Kalbfleisch and Prentice (2002), has recently been applied to conversational settings, e.g., to predict dialogue termination or disruptions De Kock and Vlachos (2021); Maystre and Russo (2022). These works, however, focus on user- or session-level outcomes and do not address the internal consistency of LLM responses under adversarial pressure. In contrast, we model time-to-inconsistency: the first incorrect answer in a multi-turn adversarial dialogue. We combine Cox proportional hazards, Accelerated Failure Time, and Random Survival Forest models and link their behavior to semantic drift covariates, enabling a nuanced statistical characterization of error accumulation and offering novel insights into dialogue reliability dynamics previously observed only empirically.
@@ 3.2 Predictive Feature Engineering
We encode each user prompt with a sentence-transformer model Reimers and Gurevych (2019):
@@ 3.3 Survival Modeling Framework
As a flexible non-parametric baseline, we employ Random Survival Forests (RSF) Ishwaran et al. (2008), which fit an ensemble of survival trees on bootstrap samples. At each split, candidate covariates are sampled at random and chosen to maximize a survival impurity reduction (log-rank statistic). Each terminal node yields a Nelson–Aalen estimate of the cumulative hazard; the forest prediction for conversation $i$ is obtained by averaging cumulative hazards across trees and converting to survival probabilities. RSF can capture nonlinearities and high-order interactions between drift features and model identity without explicit parametric assumptions.
@@ 4.1 Data
We conduct our study on the MT-Consistency Li et al. (2025a), which systematically probes LLM consistency under adversarial multi-turn interactions. Each conversation is built from a base question followed by up to 8 adversarial follow-ups; we adopt this 8-turn horizon in all experiments.
@@ 7 Conclusion
Acknowledgments
We acknowledge the fellowship support provided to Y.L. by the Center for Machine Learning and Health at Carnegie Mellon University.
This research was also funded by the National Institute of Standards and Technology under Federal Award ID Number 60NANB24D231 and Carnegie Mellon University AI Measurement Science and Engineering Center (AIMSEC).
@@ B.1 Complete Subject-to-Cluster Mappings
computer science, computer security,
@@
General Knowledge (2 subjects) truthful qa, common sense